OpenStudy (anonymous):

Solve for x. |x - 2| + 10 = 12 Solve for x. |x| = -5

OpenStudy (anonymous):

1st one: x=0 and x=4 2nd one: Not possible

OpenStudy (anonymous):

2. The absolute value of a number is always positive so getting -5 is not possible

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

cole you understood?

OpenStudy (anonymous):

\[|x-2|=12-10\] \[|x-2|=2\] \[x-2=2 \] or \[x-2=-2\]

OpenStudy (anonymous):

Start by subtracting 10 from each side, just like you would have done with a standard equation

OpenStudy (anonymous):

|x-2| can be either 2 och -2 since the negative sign is just positive inside the absolute value

OpenStudy (anonymous):

The answer is then: x= 4 and x=0

OpenStudy (anonymous):

okay thanks for showing me

OpenStudy (anonymous):

No problem